What are the vertices and the foci of this conic.
9x^2+ 4y^2 + 64y = -220
I basically need the most help with how to solve it, not just simply the answer please.
Thank you

Answer by lwsshak3(6742)

What are the vertices and the foci of this conic.
9x^2+ 4y^2 + 64y = -220
I basically need the most help with how to solve it, not just simply the answer please.
**
9x^2+ 4y^2 + 64y = -220
complete the square
9x^2+4(y^2+16y+64)=-220+256=36
9x^2+4(y+8)^2=36
divide by 36
x^2/4+(y+8)^2/9=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1,a>b, with (h,k) being the (x,y) coordinates of the center.
..
For given problem:
Center:(0,-8)
a^2=9
a=√9=3
vertices: (0,-8±a)=(0,-8±3)=(0,-11) and (0,-5)
..
b^2=4
..
c^2=a^2-b^2=9-4=5
c=√5≈2.24
Foci: (0,-8±c)=(0,-8±√5)=(0,-10.24) and (0,-5.76)